# Automatic Depression Assessment using Machine Learning: A Comprehensive Survey

### v1

@@ 2.2.6. Other depression-related human indicators
Figure 2. Illustration of common human behavioural biomarkers associated with depression.
3. Survey method
@@
This section presents our strategy to identify the papers included in the survey by reporting the procedure, search query, inclusion/exclusion criteria, selection process and publication analysis, while a structured four-phase flow figure for reporting items of our survey guided by (Moher et al., 2010), example search queries and the terminologies used in this survey are provided in the Supplementary Material.
@@
Procedure: we follow (Nightingale, 2009) and (Moher et al., 2010) to systematically review previous ADA publications, ensuring the quality and replicability of this survey, as well as the assessment of each study’s strengths and weaknesses.
@@
Eligibility Criteria: While it is impossible to exhaustively summarise all ADA approaches in this survey, papers were included only if: (i) publication dates range between 2012 and 2024; (2) they provide adequate description of ADA algorithms, with clear experimental settings and results; (3) their title, abstract or keywords contain at least one keyword describing ADA-related technology and one keyword from the Search Query Keywords; (4) they propose new ADA solutions rather than only applying them; and (5) they published in English. Meanwhile, papers were excluded if: (1) they did not involve any ML algorithm for inferring depression status; and (2) they were not under peer-review process or come from sources that may not meet credibility standards.
Selection Process: All candidate publications were initially searched based on their titles and abstracts, followed by a full-text review to ensure compliance with eligibility criteria. To establish consensus, we assessed a random sample of three papers for each modality, with full agreement on details of inclusion and exclusion criteria. The remaining papers were randomly assigned to at least two authors for independent screening, where ambiguous cases were resolved through discussion. This way, a total of 1679 papers were retrieved, with 183 were ultimately included. The full list of the included papers are provided in the Supplementary Material.
Publication analysis: We defined a set of variables to understand key information from each ADA paper using an analytical approach, including numerical variables such as depression classification accuracy, the numbers of participants and clips, as well as categorical variables such as modality, methodology type, predictor and fusion strategy (summarised in Table 1, with the full list of variables provided in the Supplementary Material). A descriptive statistics description is employed to represent numerical variables, while categorical variables were derived through either a top-down or bottom-up strategy.
4. Automatic depression assessment
Table 1. Summary of the surveyed ADA approaches, where ‘C’ and ‘R’ denote the classification and regression, respectively; ‘Hybrid’ denotes approaches utilising both traditional ML and DL techniques; IAIF: Iterative Adaptive Inverse Filtering; OQNN: Open Quotient Neural Network; DAE: Denoising Auto-Encoder denotes; EOH: Edge Orientation Histogram; ADTP: Audio Delta Ternary Patterns; ReHo: Regional Homogeneity; FALFF: Fractional Amplitude of Low Frequency Fluctuations.
@@ 4. Automatic depression assessment
EEG C ML GA, LDA, EMD, CBEM, GSW SVM, KNN, RF, NB, DT (Shen et al., 2023), (Mohammadi et al., 2015), (Zhu et al., 2020), (Shen et al., 2020)
DL CNN, CNN+LSTM/GRU, MLP, GNN, Transformer MLP (Dang et al., 2020), (Wan et al., 2020), (Sharma et al., 2023), (Shao et al., 2024b), (Wang et al., 2024b), (Lu et al., 2024), (Shen et al., 2023), (Ay et al., 2019), (Song et al., 2022c), (Tigga and Garg, 2022), (Tian et al., 2024), (Ying et al., 2024)
Hybrid GNN SVM, MLP (Zhang et al., 2024b), (Li et al., 2023), (Wang et al., 2021a)
MRI C ML ReHo, FALFF, ICA, RFE, LASSO SVM, LDA, GPC, KNN, DT, LR, RF, GNB, BNB, DT, TVF, SVF (Mousavian et al., 2021), (Gao et al., 2018), (Winter et al., 2024)
DL CNN, GNN MLP (Chen et al., 2021b), (Gao et al., 2023), (Dai et al., 2023), (Zhu et al., 2023), (Yao et al., 2020), (Pominova et al., 2018), (Lin et al., 2023), (Jun et al., 2020), (Ktena et al., 2018), (Fang et al., 2023b), (Qin et al., 2022), (Jaiswal et al., 2019), (Li et al., 2024a)
Hybrid CNN, LSTM BernoulliNB, MLP (Mousavian et al., 2020)
C & R DL GNN MLP (Kong et al., 2021)
Visual C ML PCA + LDA, AAM, LBPH, LBP-TOP, HOG, RBM, GMM KNN, SVM, GMM, RF, NBC, TF, SG, PSD (Gahalawat et al., 2023), (Girard et al., 2013), (Anis et al., 2018), (Joshi et al., 2013b), (Girard et al., 2014), (Harati et al., 2019), (Yang et al., 2023a), (Al-gawwam and Benaissa, 2018), (Yang et al., 2022), (Alghowinem et al., 2015), (Wang et al., 2020a), (Wang et al., 2021b), (Joshi et al., 2012), (Joshi et al., 2013a), (Lu et al., 2022)
DL MLP, CNN, Transformer MLP (Li et al., 2024b), (Li et al., 2024c), (Xie et al., 2021), (Guo et al., 2022), (Niu et al., 2024b)
Hybrid LBP-TOP, MRLBP-TOP, CNN, CNN-LSTM DT, SVM, KNN, DT, MLP (He et al., 2018), (Shao et al., 2021), (Yu et al., 2022), (Liu et al., 2024)
R ML LPQ, LGBP-TOP, MRLBP-TOP, LPQ-TOP SVR (Tao et al., 2019), (Du et al., 2019), (Dhall and Goecke, 2015), (Wen et al., 2015), (Song et al., 2024)
DL MLP, CNN, CNN+RNN, GRU, CNN+Transformer, GNN MLP (de Melo et al., 2021), (Xu et al., 2024), (Pan et al., 2023), (Ahmad et al., 2021), (Shang et al., 2021), (Song et al., 2022a), (He et al., 2022a), (Niu et al., 2024a), (Jaiswal et al., 2019), (Liu et al., 2022), (Zhao et al., 2023), (Li et al., 2020), (Zhou et al., 2018), (Uddin et al., 2020), (Pan et al., 2024b), (Al Jazaery and Guo, 2018), (Chen et al., 2021a), (Zhu et al., 2017), (Zhang et al., 2023a), (He et al., 2024), (de Melo et al., 2019), (Moreno et al., 2023), (Chen et al., 2022), (Pan et al., 2024a), (Wang et al., 2024a), (Li et al., 2024b)
C & R DL Transformer MLP (Zhang et al., 2023b)
Hybrid CNN LightGBM, SVR, MLP (Song et al., 2018), (Islam and Bae, 2024)
Audio C ML MFCC, CLP QDA, LR, GMM, MLC (Mundt et al., 2012), (Moore II et al., 2007), (Ozdas et al., 2004)
DL CNN, CNN+LSTM MLP (Vázquez-Romero and Gallardo-Antolín, 2020), (Ma et al., 2016)
Hybrid COVAREP, CNN, LSTM RF, DT, MLP (Zhang et al., 2021), (Wu et al., 2024)
R ML RF, LR (Sun et al., 2017)
DL CNN, LSTM, Transformer MLP (He and Cao, 2018), (Al Hanai et al., 2018), (Amiriparian et al., 2017), (Zhao et al., 2019)
Hybrid MFCC, ADTP, rPPG, LGBP-TOP, HRV, CNN, CNN-LSTM RF, MLP (Casado et al., 2023), (Fu et al., 2022)
C & R ML SVM, RF (Valstar et al., 2016)
DL CNN, RNN/LSTM MLP (Othmani et al., 2021), (Rejaibi et al., 2022), (Muzammel et al., 2020), (Zhang et al., 2021)
Langua- ge/text C ML N-gram, LDA, LIWC LR, SVM, RF, AdaBoost, DT, UDL, TF-IDF (De Choudhury et al., 2013b), (De Choudhury et al., 2013a), (Nguyen et al., 2014), (Tsugawa et al., 2015), (Shen et al., 2017), (Tadesse et al., 2019)
DL CNN, GRU, Transformer MLP (Trotzek et al., 2018), (Pérez-Toro et al., 2024), (Khowaja et al., 2024)
R ML Unigrams, LIWC, LDA LR, SVR (Resnik et al., 2015), (Resnik et al., 2013)
Multi- modal C ML HMMs, AMMs, FACS KNN, SVM, DT, NB (Pampouchidou et al., 2016), (Cohn et al., 2009), (Pampouchidou et al., 2015), (Nasir et al., 2016), (Senoussaoui et al., 2014), (Alghowinem et al., 2016), (Ning et al., 2024)
DL CNN, LSTM/GRU, Transformer MLP (Yoon et al., 2022), (Shen et al., 2022), (Ye et al., 2021), (Cheong et al., 2024b), (Wei et al., 2022), (Zheng et al., 2023), (Shao et al., 2024a), (Tao et al., 2024), (Ye et al., 2024), (Jung et al., 2024)
Hybrid Mfcc, DAE, COAVREP, FACS SVM, RF, LR, MLP (Lin et al., 2021), (Dibeklioğlu et al., 2018), (Cheong et al., 2024c)
R ML LBP, EOH, LPQ, MHH, COVAREP, LIWC RF, PLS, SVR, RF (Gong and Poellabauer, 2017), (Valstar et al., 2013), (Valstar et al., 2014), (Ringeval et al., 2017), (Gratch et al., 2014), (Jan et al., 2014), (Gupta et al., 2014), (Meng et al., 2013), (Kächele et al., 2014)
DL CNN, LSTM/GRU, CNN+LSTM, Transformer, GNN MLP (He et al., 2021), (Niu et al., 2020), (Rodrigues Makiuchi et al., 2019), (Ringeval et al., 2019), (He et al., 2022c), (Kaya et al., 2019), (Hong et al., 2022), (Uddin et al., 2022), (Hu et al., 2024), (Sun et al., 2022b), (Sun et al., 2022a), (Yin et al., 2019), (Fang et al., 2023a), (Zhang, [n. d.]), (Fan et al., 2024), (Gimeno-Gómez et al., 2024)
Hybrid CNN, GNN SVR, PLS, MLP (Jan et al., 2017), (Shen et al., 2024)
C & R ML FAUs, dMFCCs, COVAREP, IAIF, OQNN SVM, DT (Yang et al., 2016), (Williamson et al., 2016), (Scherer et al., 2013a)
DL CNN+LSTM, GNN, Transformer MLP (Muzammel et al., 2021), (Zhang et al., 2024a), (Chen et al., 2023), (Lam et al., 2019), (Zhang et al., 2024c)
Hybrid CNN SVM, RF, MLP (Yang et al., 2017), (Yang et al., 2018), (Haque et al., 2018)
@@
4.1. EEG-based automatic depression assessment
4.1.1. Challenges.
@@ 4.1. EEG-based automatic depression assessment
While a significant depression-related EEG biomarker is its spectral features, whose principal bands can be divided into five parts: Delta, Theta, Alpha, Beta, and Gamma describing different physiological meanings, a key challenge unique to EEG-based systems is determining which bands to use for ADA. Hosseinifard et al. (Hosseinifard et al., 2013) found that the combination of alpha and theta bands is a discriminative depression feature. When logistic regression (LR) is applied to process such features, 90% depression classification accuracy is achieved for 90 test subjects. Other challenges unique to EEG-based ADA, including: (i) how to properly model the spatial information derived from multiple EEG channels; and (ii) the varying importance of different channels (Shen et al., 2023; Zhang et al., 2024b). In this sense, Shen et al. (Shen et al., 2020) modify kernel target alignment (KTA) strategy to optimally choose a subset of important EEG channels (addressing problem (ii)) for reducing the information redundancy, which are then processed by a CNN for EEG-based ADA. Their later work (Shen et al., 2023) develops an adaptive channel fusion framework to model the spatial information derived from multi-channel EEG signals (addressing problem (i)), where two improved focal loss functions are employed to upweight the loss values for the hard examples. In addition, recent studies (Zhang et al., 2024b; Li et al., 2023; Wang et al., 2021a) frequently treat EEG signals as graphs to address these three problems, where each EEG channel is represented as a node, with edges are defined to capture relationships between connected EEG channels. Subsequently, GNN with an attention module (Zhang et al., 2024b), or Multi-scale Adaptive Spatial-Temporal GNN (Lu et al., 2024) have been proposed to weight and jointly process spatial cues of multiple EEG channels. We further discuss other EEG-based ADA methods in the following subsection.
4.1.2. Other EEG-based ADA systems

### v2

@@ 2.2.1 Brain cues
Figure 2: Illustration of common human behavioural biomarkers associated with depression.
2.2.2 Physiological cues
Depression is also associated with abnormal physiological states, as changes in the brain affects on various body systems. For instance, fluctuations in serotonin that are related to depression frequently cause depressed individuals suffering from sleep problems [9] and lower sex drive [100]. The state of ‘accelerated aging’ [14] is another typical depression biomarker, as depression causes comparable changes on human brain’s plasticity and synaptic functions [101] that are similar to those occurring with aging, where depressed individuals usually suffer from 4–6 years of accelerated aging compared to normal people [102].
2.2.3 Physical indicators
Meanwhile, Katon et al. [103] and Lin et al. [104] concluded that depression leads individuals to experience decreased heart rate variability, increased adhesiveness of platelets, negative health habits such as smoking and over-eating, and low circadian rhythmicity in their skin body temperatures [105]. Appetite change is another physical indicator of depression, where greater appetite change tends to indicate greater depression severity [8]. As a result, human body shape and weight changes are important physical indicators to indicate depression [106], e.g., the obesity [15], where a reciprocal and reinforcing cycle of maladaptive physiological adjustments can establish an interconnection between the depression and obesity [16]. As discussed in Sec. 2.1, sickness-like behaviours, cognitive dysfunction, and anhedonia [107] are also typical depression-related cues. For example, there is a bidirectional relationships linking depression to the progression of cancer [108] and linking the persistent inflammatory processes to stresses in the manifestation of depressive symptoms [109].
2.2.4 Non-verbal expressive behaviours
Depressed individuals typically express different non-verbal expressive behaviours, which facilitates existing audio-visual ADA approaches.
Gesture and body movements: Depressed individuals typically show physical activity known as psychomotor retardation [17], represented by different gait from normal people [19] such as slumped, hunched, or rigid posture that restricts body movements [110]), impaired gesture movements (e.g., gesture deficits and degraded intransitive behaviours) [111], fewer gestures during interpersonal interactions [18], as well as downwards head-tilting and frequent self-adaptor, self-touch or hand-over-face gestures [112, 113, 67].
Facial behaviours: A typical facial depression biomarker is the reduced or even diminished facial expressions [17, 114] and head movements [115], especially significantly reduced expressions of positive emotions. However, contradicting conclusions have been reached in terms of negative facial expressions, where some studies claimed that depressed individuals experience increased negative expressions [116] while others shown that even negative facial expressions are less expressed by them [114]. From a more fine-grained perspective, Cohn et al. [117] and Girard et al. [118] conducted pioneering investigations regarding the depression-related facial AU patterns, both of which show that participants suffering from high depression severity display less affiliative facial expressions (AU 12 and AU 15) but more non-affiliative facial expressions (AU 14). Also, an early study [119] suggested that depressed individuals tend to avoid eye contact, and show more of nonspecific gaze and less mutual gaze, indicating lack of interest in communication.
Audio behaviours: Non-verbal audio behaviours are also crucial depression biomarkers [120], as depression-related cognitive impairments would affect human working memory that controls the articulatory system and speech-based information storage [121]. Specifically, depressed individuals typically exhibit monotonous speech [17] characterised by frequent pauses with a ‘muffled’ voice [122], reduction in articulatory coordination [123] and verbal fluency [124], abnormal pitch variability, vocalization time, fundamental frequency and formants [125, 21], long pause times [125], as well as an overall decrease in speech rate/speed [23]. An early study summarised depressive voices as ‘patients speak in a low voice, slowly, hesitatingly, monotonously, sometimes stuttering, whispering, try several times before they bring out a word, become mute in the middle of a sentence’ [126]. Besides, Hollien et al. [23] found that depressed individuals are more likely to experience reduced speaking intensity, pitch range and intonation, as well as a lack of linguistic stress, while Breznitz et al. [127] revealing that narrower frequencies are expressed by depressed females’ speeches compared to health females.
2.2.5 Verbal language behaviours
Language-based communication is a primary mean for people to exchange their ideas and intentions, and thus its styles are strongly associated with their depression status [21], despite that there exists cross-cultural differences in depression-related verbal expressions [128] (e.g., specific negative emotions or phrases are stronger depression indicators in Chinese social platforms than in other western platforms [129]). Among them, the most recognised verbal depression biomarkers are the increase usage of the first person singular (i.e., ‘I’, ‘me’, and ‘mine’) [130], swear words and depression terms, as well as the decrease usage of 3rd person pronoun (e.g., ‘She’, ‘He’). These are evidenced by the studies conducted on suicidal poets [131], student essays [132], Twitter [133] and a Chinese social platform - Sina [129]. A survey [134] further revealed that individuals with depression and suicidal tendencies have an increased usage of self-referential words such as ’I’, ’me’, and ’mine’. They additionally found that depressed individuals have a general decrease usage of 3rd person pronoun (e.g., ’She’, ’He’), and a increase over time for swear word use and frequency of depression terms. Other studies [135] further found that tentative words (e.g., maybe, perhaps), negations (e.g., no, never), adverbs, and conjunctions are more commonly used by depressive individuals, with less usage of verbs, while postpartum depression mothers tending to show more non-fluency in their social platform posts. From another perspective, Rude et al. [132] found that depressed individuals used more negatively valenced words, while Cummins et al. [136] summarised that they are more likely to express dull, misery and “lifeless” descriptors. Besides, the sentences expressed by depressed patients are usually fragmented or shortened, reflecting their reduced coherence and cognitive difficulty [137].
2.2.6 Other depression-related human indicators
In addition to the behaviours discussed above, depressed patients also show various cognitive deficits compared to non-depressed ones, including varied exploratory behaviours [138], impaired recognition of basic emotional facial expressions [4], longer reaction time, a decreased psychomotor ability [139] and diminished capability in acquiring new information (but retaining the capability in gathering information) [139]. Memory abnormality is a key sign reflecting depression, as depression causes the suppression of hippocampal neurogenesis, inhibition of dopamine neurons, and increased sensitivity of the amygdala [140]. As evidenced by [141], depressed individuals tend to exhibit difficulties in recalling information, i.e., they frequently experience enhanced memory for negative content but poorer memory for positive material [142]. With such cognitive and memory deficits and bias, multiple human cognitive capabilities are further negatively impacted, i.e., depressed individuals frequently show abnormality or lower performances in decision-making [138], problem-solving [143], planning [144] behaviours, as well as spending time with similarly depressed others and pair-wise interactions rather than group interactions [145].
3 Survey Method
@@
This section presents our strategy to identify the papers included in the survey by reporting the procedure, search query, inclusion/exclusion criteria, selection process and publication analysis, while a structured four-phase flow figure for reporting items of our survey guided by [146] is illustrated in Fig. 3. Example search queries and the terminologies used in this survey are provided in the Supplementary Material.
Figure 3: PRIMA Schema for our systematic review.
@@
Procedure: we follow [147] and [146] to systematically review previous ADA publications, ensuring the quality and replicability of this survey, as well as the assessment of each study’s strengths and weaknesses.
@@
Eligibility Criteria: While it is impossible to exhaustively summarise all ADA approaches in this survey, papers were included only if: (i) publication dates range between 2012 and 2024; (2) they provide adequate description of ADA algorithms, with clear experimental settings and results; (3) their title, abstract or keywords contain at least one keyword describing ADA-related technology and one keyword from the Search Query Keywords; (4) they propose new ADA solutions rather than only applying them; and (5) they published in English. Meanwhile, papers were excluded if: (1) they did not involve any ML algorithm for inferring depression; and (2) they were not under peer-review process or come from sources that may not meet credibility standards.
Selection Process: All candidate publications were initially searched based on their titles and abstracts, followed by a full-text review to ensure compliance with eligibility criteria. To establish consensus, we assessed a random sample of three papers for each modality, with full agreement on details of inclusion and exclusion criteria. The remaining papers were randomly assigned to at least two authors for independent screening, where ambiguous cases were resolved through discussion. This way, a total of 1679 papers were retrieved, with 183 were ultimately included (the full list is provided in the Supplementary Material).
Publication analysis: We defined a set of variables to understand key information from each ADA paper using an analytical approach, including numerical variables such as depression classification accuracy, the numbers of participants and clips, as well as categorical variables such as modality, methodology type, predictor and fusion strategy (summarised in Table I, with the full list of variables provided in the Supplementary Material). A descriptive statistics description is employed to represent numerical variables, while categorical variables were derived through either a top-down or bottom-up strategy.
4 Automatic Depression Assessment
TABLE I: Summary of the surveyed ADA approaches, where ‘C’ and ‘R’ denote the classification and regression, respectively; ‘Hybrid’ denotes approaches utilising both traditional ML and DL techniques; IAIF: Iterative Adaptive Inverse Filtering; OQNN: Open Quotient Neural Network; DAE: Denoising Auto-Encoder; EOH: Edge Orientation Histogram; ADTP: Audio Delta Ternary Patterns; ReHo: Regional Homogeneity; FALFF: Fractional Amplitude of Low Frequency Fluctuations.
@@ 4 Automatic Depression Assessment
EEG C ML GA, LDA, EMD, CBEM, GSW SVM, KNN, RF, NB, DT [45], [148], [149], [150]
DL CNN, CNN+LSTM/GRU, MLP, GNN, Transformer MLP [151], [152], [52], [53], [54], [55], [45], [153], [154], [155], [156], [157]
Hybrid GNN SVM, MLP [44], [56], [158]
MRI C ML ReHo, FALFF, ICA, RFE, LASSO SVM, LDA, GPC, KNN, DT, LR, RF, GNB, BNB, DT, TVF, SVF [48], [71], [159]
DL CNN, GNN MLP [46], [47], [57], [58], [160], [161], [162], [163], [164], [165], [166], [167], [168]
Hybrid CNN, LSTM BernoulliNB, MLP [169]
C & R DL GNN MLP [170]
Visual C ML PCA + LDA, AAM, LBPH, LBP-TOP, HOG, RBM, GMM KNN, SVM, GMM, RF, NBC, TF, SG, PSD [33], [39], [171], [115], [118], [172], [173], [174], [175], [176], [177], [178], [179], [180], [181]
DL MLP, CNN, Transformer MLP [182], [183], [184], [185], [186]
Hybrid LBP-TOP, MRLBP-TOP, CNN, CNN-LSTM DT, SVM, KNN, DT, MLP [187], [188], [189], [190]
R ML LPQ, LGBP-TOP, MRLBP-TOP, LPQ-TOP SVR [191], [192], [193], [194], [195]
DL MLP, CNN, CNN+RNN, GRU, CNN+Transformer, GNN MLP [43], [28], [35], [40], [41], [42], [196], [197], [167], [198], [199], [200], [201], [202], [203], [204], [205], [206], [207], [208], [209], [210], [211], [212], [38], [182]
C & R DL Transformer MLP [213]
Hybrid CNN LightGBM, SVR, MLP [214], [215]
Audio C ML MFCC, CLP QDA, LR, GMM, MLC [125], [216], [217]
DL CNN, CNN+LSTM MLP [218], [219]
Hybrid COVAREP, CNN, LSTM RF, DT, MLP [220], [221]
R ML RF, LR [222]
DL CNN, LSTM, Transformer MLP [223], [224], [225], [226]
Hybrid MFCC, ADTP, rPPG, LGBP-TOP, HRV, CNN, CNN-LSTM RF, MLP [227], [228]
C & R ML SVM, RF [61]
DL CNN, RNN/LSTM MLP [229], [230], [231], [220]
Langua- ge/text C ML N-gram, LDA, LIWC LR, SVM, RF, AdaBoost, DT, UDL, TF-IDF [133], [135], [232], [128], [233], [234]
DL CNN, GRU, Transformer MLP [235], [236], [30]
R ML Unigrams, LIWC, LDA LR, SVR [137], [237]
Multi- modal C ML HMMs, AMMs, FACS KNN, SVM, DT, NB [25], [117], [238], [239], [240], [241], [242]
DL CNN, LSTM/GRU, Transformer MLP [68], [69], [243], [244], [245], [246], [247], [248], [31], [27]
Hybrid Mfcc, DAE, COAVREP, FACS SVM, RF, LR, MLP [67], [66], [249]
R ML LBP, EOH, LPQ, MHH, COVAREP, LIWC RF, PLS, SVR, RF [36], [59], [60], [62], [64], [250], [251], [252], [253]
DL CNN, LSTM/GRU, CNN+LSTM, Transformer, GNN MLP [254], [29], [37], [63], [255], [256], [257], [258], [259], [260], [261], [262], [32], [263], [264], [20]
Hybrid CNN, GNN SVR, PLS, MLP [265], [266]
C & R ML FAUs, dMFCCs, COVAREP, IAIF, OQNN SVM, DT [267], [268], [269]
DL CNN+LSTM, GNN, Transformer MLP [26], [270], [271], [272], [273]
Hybrid CNN SVM, RF, MLP [274], [275], [276]
@@
4.1 EEG-based automatic depression assessment
Non-invasive electroencephalography (EEG) is a popular sensor for monitoring human brain electrical activities that are related to human depression status. Thus, EEG-based ADA solutions have been extensively developed.
4.1.1 Challenges and solutions
@@
While a significant depression-related EEG biomarker is its spectral features, whose principal bands can be divided into five parts: Delta, Theta, Alpha, Beta, and Gamma describing different physiological meanings, a key challenge unique to EEG-based systems is determining which bands to use for ADA. Hosseinifard et al. [50] found that the combination of alpha and theta bands is a discriminative depression feature. When logistic regression (LR) is applied to process such features, 90% depression classification accuracy is achieved for 90 test subjects. Other challenges unique to EEG-based ADA, including: (i) how to properly model the spatial information derived from multiple EEG channels; and (ii) the varying importance of different channels [45, 44]. In this sense, Shen et al. [150] modify kernel target alignment (KTA) strategy to optimally choose a subset of important EEG channels (addressing problem (ii)) for reducing the information redundancy, which are then processed by a CNN for EEG-based ADA. Their later work [45] develops an adaptive channel fusion framework to model the spatial information derived from multi-channel EEG signals (addressing problem (i)), where two improved focal loss functions are employed to upweight the loss values for the hard examples. In addition, recent studies [44, 56, 158] frequently treat EEG signals as graphs to address these three problems, where each EEG channel is represented as a node, with edges are defined to capture relationships between connected EEG channels. Subsequently, GNN with an attention module [44], or Multi-scale Adaptive Spatial-Temporal GNN [55] have been proposed to weight and jointly process spatial cues of multiple EEG channels. We further discuss other EEG-based ADA methods in the following subsection.
4.1.2 Other EEG-based ADA systems